### v3

# Similarity-Dissimilarity Loss for Multi-label Supervised Contrastive Learning

Abstract
Supervised contrastive learning has achieved remarkable success by leveraging label information; however, determining positive samples in multi-label scenarios remains a critical challenge. In multi-label supervised contrastive learning (MSCL), relations among multi-label samples are not yet fully defined, leading to ambiguity in identifying positive samples and formulating contrastive loss functions to construct the representation space. To address these challenges, we: (i) first define five distinct multi-label relations in MSCL to systematically identify positive samples, (ii) introduce a novel Similarity-Dissimilarity Loss that dynamically re-weights samples through computing the similarity and dissimilarity factors between positive samples and given anchors based on multi-label relations, and (iii) further provide theoretical grounded proof for our method through rigorous mathematical analysis that supports the formulation and effectiveness of the proposed loss function. We conduct the experiments across both image and text modalities, and extend the evaluation to medical domain. The results demonstrate that our method consistently outperforms baselines in a comprehensive evaluation, confirming its effectiveness and robustness. Code is available at: https://github.com/guangminghuang/similarity-dissimilarity-loss.
Index Terms:
Multi-label Supervised contrastive learning (MSCL), multi-label classification, international classification of diseases (ICD).
I Introduction
Multi-label classification presents significant challenges due to its inherent label correlations, extreme and sparse label spaces, and long-tailed distributions. For instance, in the International Classification of Diseases (ICD) [1, 2], the presence of one label (e.g., ”Pneumococcal pneumonia”) may increase the probability of co-occurring labels (e.g., ”fever” or ”cough”). Furthermore, multi-label datasets frequently exhibit long-tailed distributions, where a small subset of labels occurs with high frequency while the majority appear rarely. This imbalance typically results in models that perform adequately on common labels but underperform on infrequent ones [3, 4]. Additionally, the number of potential label combinations increases exponentially with the number of labels, resulting in heightened computational complexity and substantial memory requirements.
Supervised contrastive learning effectively utilizes label information to yield promising results in single-label scenarios [5]. However, identifying positive samples in multi-label supervised contrastive learning (MSCL) remains a challenge. For example, consider a set of images containing cats and puppies, wherein an anchor image depicts a cat; in the single-label paradigm, positive and negative instances can be unambiguously delineated based on their corresponding taxonomic annotations. Conversely, MSCL introduces inherent classification ambiguity when determining whether an image containing both cats and puppies should be designated as a positive or negative sample in relation to the anchor.
A critical question arises: Should a sample be considered positive when its label set partially overlaps with or exactly matches that of the anchor? Currently, three principal strategies exist for identifying positive samples in multi-label scenarios [6]: (i) ALL considers only samples with an exactly matching label set as positive; (ii) ANY identifies samples with any overlapping class with the anchor as positive, and (iii) MulSupCon [6] conceptually aligns with the ANY approach but treats each label independently, thereby generating multiple distinct positive sets for individual anchor samples.
However, these methods have inherent limitations, since previous research has overlooked the complicated multi-label relations among samples in MSCL. As illustrated in Figure 1, we introduce five distinct set relations among samples to facilitate a more comprehensive identification of positive sets. The ALL strategy exclusively considers relation $R2$ while disregarding the potential contributions of $R3$, $R4$ and $R5$. Furthermore, long-tailed distributions, when tail samples serve as anchors, the ALL strategy’s requirement for exact label matches significantly impedes these tail anchors from identifying adequate positive samples within a limited batch size, potentially degenerating the method to unsupervised contrastive learning in extreme scenarios [7, 8, 3]. Conversely, both ANY and MulSupCon approaches treat relations $R2$, $R3$, $R4$, and $R5$ identically with equivalent weights in contrastive loss functions, which constitutes a suboptimal approach given the inherent differences among these relations. A detailed mathematical analysis of these three methods is presented in Section II.
Figure 1: Five distinct relations between samples and a given anchor. $\Omega$ denotes a universe that contains all label entities. Here is an example with five different relations between sample $p$ and anchor $i$, where the labels are represented as one-hot vectors.
To address these issues, we define multi-label relations and introduce a novel contrastive loss function. Our contributions are summarized as follows:
  1. 1.
    To the best of our knowledge, we are the first to define multi-label relations in MSCL, which facilitates the identification of complex relations in multi-label scenarios.
  2. 2.
    We introduce similarity and dissimilarity concepts in multi-label scenarios and propose a novel contrastive loss function, termed Similarity-Dissimilarity Loss, which dynamically re-weights based on the computed similarity and dissimilarity factors between positive samples and anchors, guided by multi-label relations.
  3. 3.
    We establish the theoretical foundations of our approach through rigorous mathematical analysis, demonstrating both the formal derivation, and the upper and lower bounds of the weighting factor.
  4. 4.
    We conduct the experiments across both image and text modalities, and extend the evaluation to medical domain. The results demonstrate that our method consistently outperforms baselines in a comprehensive evaluation, confirming its effectiveness and robustness.
II Methods
In this section, we establish the preliminary notation and adhere to the conventions established in [5] to maintain consistency throughout our analysis. Subsequently, we examine the limitations of the ALL, ANY, and MulSupCon strategies and their corresponding loss functions. We then introduce our formulation of multi-label relations and present the Similarity-Dissimilarity Loss for MSCL. Furthermore, we provide a rigorous mathematical analysis to establish the theoretical foundations of the proposed methodology.
II-A Preliminaries
Given a batch of $N$ randomly sample/label pairs, $\{(\boldsymbol{x}_{i},\boldsymbol{y}_{i})\}_{i=1,\dots,N}$, where $\boldsymbol{x}_{i}$ denotes the $i$-th sample and $\boldsymbol{y}_{i}$ its corresponding labels. Here, $\boldsymbol{y}_{i}=\{y_{i}^{(l)}\}_{l=1,\dots,L}$ represents the multi-labels of sample $i$, where $y_{i}^{(l)}$ denotes the $l$-th label of sample $i$ and $L$ is the total number of labels for sample $i$. After data augmentation, the training batch consists of $2N$ pairs, $\{\tilde{\boldsymbol{x}}_{j},\tilde{\boldsymbol{y}}_{j}\}_{j=1,\dots,2N}$, where $\tilde{\boldsymbol{x}}_{2i}$ and $\tilde{\boldsymbol{x}}_{2i-1}$ are two random augmentations of $\boldsymbol{x}_{i}$ ($i=1,\dots,N$) and $\tilde{\boldsymbol{y}}_{2i-1}=\tilde{\boldsymbol{y}}_{2i}=\boldsymbol{y}_{i}$. For brevity, we refer to this collection of $2N$ augmented samples as a ”batch” [5].
II-B Multi-label Supervised Contrastive Loss
In MSCL, the formulation of supervised contrastive loss varies depending on the strategies employed for determining positive samples relative to a given anchor. Let $i\in\mathcal{I}=\{1,\dots,2N\}$ denote the index of an arbitrary augmented sample. For the ALL strategy, the positive set is defined as follows:
$\mathcal{P}(i)=\{p\in\mathcal{A}(i)|\forall p,\tilde{\boldsymbol{y}}_{p}=\tilde{\boldsymbol{y}}_{i}\}$ (1)
where $\mathcal{A}(i)\equiv I\setminus\{i\}$  [^1] In contrastive learning, sample $i$ is the anchor and is supposed to be excluded out of positive sets..
Subsequently, the positive set for the ANY strategy is defined as follows:
$\mathcal{P}(i)=\{p\in\mathcal{A}(i)|\forall p,\tilde{\boldsymbol{y}}_{p}\cap\tilde{\boldsymbol{y}}_{i}\neq\varnothing\}$ (2)
In MSCL, the form of contrastive loss function for ALL and ANY is identical. For each anchor $i$, the loss function is formulated as follows:
$\mathcal{L}_{i}=\frac{-1}{|\mathcal{P}(i)|}\sum_{p\in\mathcal{P}(i)}\log\frac{\exp(\boldsymbol{z}_{i}\cdot\boldsymbol{z}_{p}/\tau)}{\sum_{a\in\mathcal{A}(i)}\exp(\boldsymbol{z}_{i}\cdot\boldsymbol{z}_{a}/\tau)}$ (3)
Here, $\tau\in\mathbb{R}^{+}$ represents a positive scalar temperature parameter [7], while $\boldsymbol{z}_{k}=Proj(Enc(\tilde{\boldsymbol{x}}_{k}))\in\mathbb{R}^{D_{P}}$ denotes the projected encoded representation [5].
For a given batch of samples, the loss function is formulated as:
$\mathcal{L}=\sum_{i\in I}\mathcal{L}_{i}$ (4)
Zhang et la [6] propose an approach that considers each label $\tilde{y}_{i}^{(l)}$ independently, forming multiple positive sets for a given anchor sample $i$. For each label $\tilde{y}_{i}^{(l)}\in\tilde{\boldsymbol{y}}_{i}$, the positive set for the MulSupCon is defined as:
$\mathcal{P}(i)=\{p\in\mathcal{A}(i)|\forall p,\tilde{y}_{p}^{(l)}\in\tilde{\boldsymbol{y}}_{i}\}$ (5)
For each anchor $i$, the multi-label supervised contrastive loss for MulSupCon is represented as follows [6]:
$\mathcal{L}_{i}^{mul}=\sum_{\tilde{y}_{p}^{(l)}\in\tilde{\boldsymbol{y}}_{i}}\frac{-1}{|\mathcal{P}(i)|}\sum_{p\in\mathcal{P}(i)}\log\frac{\exp(\boldsymbol{z}_{i}\cdot\boldsymbol{z}_{p}/\tau)}{\sum_{a\in\mathcal{A}(i)}\exp(\boldsymbol{z}_{i}\cdot\boldsymbol{z}_{a}/\tau)}$ (6)
For a given batch of samples, the loss function is formulated as:
$\mathcal{L}^{mul}=\frac{1}{\sum_{i}|\tilde{\boldsymbol{y}}_{i}|}\sum_{i\in I}\mathcal{L}_{i}^{mul}$ (7)
II-C Multi-label Relations
As illustrated in Figure 1, we denote each Relation as $R$, where, e.g., $R1$ stands for Relation 1. The subscripted notation $p_{j}$ signifies that sample $p$ corresponds to the $j$-th relation.
Let $\Omega$ denote a universal set containing all possible label entities. For any anchor $i$ and sample $p$, let $\mathcal{S}$ and $\mathcal{T}$ represent their respective label sets. The five fundamental multi-label relations are defined as follows:
$R1$ $:\mathcal{S}\cap\mathcal{T}=\varnothing$ (8)
$R2$ $:\mathcal{S}=\mathcal{T}$ (9)
$R3$ $:\mathcal{S}\cap\mathcal{T}\neq\varnothing,\mathcal{S}\nsubseteq\mathcal{T},\mathcal{T}\nsubseteq\mathcal{S}$ (10)
$R4$ $:\mathcal{S}\supsetneq\mathcal{T}$ (11)
$R5$ $:\mathcal{S}\subsetneq\mathcal{T}$ (12)
Based on these relational definitions, we present a theoretical analysis of the limitations inherent in the ALL, ANY, and MulSupCon methods, illustrated via an example in Figure 1.
In the ALL method, the optimization process aims to align with the mean representation of samples sharing identical label sets [6]. As the example that is demonstrated in Figure 1, for a given anchor $i$, the positive set of ALL is:
$\mathcal{P}(i)=\{p_{2}\}$
In the ALL method, the sample $p_{j}$ in $R2$ is designated as positive sample, while those in relations $R3$, $R4$ and $R5$ are excluded from consideration. Specifically, despite their semantic similarity to anchor $i$ that those overlap labels, the feature representations of samples $p_{j}$ where $j\in{3,4,5}$ are forced away from the anchor in the embedding space, as they are treated as negative examples in the contrastive learning paradigm. Consequently, the restricted size of the positive set $|\mathcal{P}(i)|$ results in a mean representation susceptible to statistical variance. Furthermore, the ALL method may inadvertently treat semantically related samples as negative instances in certain scenarios.
Lemma 1.
(Vector Similarity Under Label Equivalence). Let $i$ be an anchor and $p$ be any sample in the feature space, where $\tilde{\boldsymbol{y}}_{i},\tilde{\boldsymbol{y}}_{p}\in\mathbb{R}^{d}$ denote their respective label vectors. If $\tilde{\boldsymbol{y}}_{p}=\tilde{\boldsymbol{y}}_{i}$, then under the contrastive learning framework [7], their corresponding projected representations $\boldsymbol{z}_{i},\boldsymbol{z}_{p}\in\mathbb{R}^{m}$ satisfy $\boldsymbol{z}_{i}\simeq\boldsymbol{z}_{p}$.
As per ANY’s definition, the positive set of the example in Figure 1 is:
$\mathcal{P}(i)=\{p_{2},p_{3},p_{4},p_{5}\}$
By applying Lemma 1, the corresponding loss terms in Eq. (3) for samples in different relations exhibit approximate equality:
$\mathcal{L}(R2)\approx\mathcal{L}(R3)\approx\mathcal{L}(R4)\approx\mathcal{L}(R5)$
It is evident that $R2$, $R3$, $R4$ and $R5$ represent fundamentally distinct relations, each characterized by different labels and semantic information. However, the ANY method fails to differentiate these subtle label hierarchies, introducing substantial semantic ambiguity. Moreover, in scenarios where samples predominantly share common classes, the averaging mechanism disproportionately emphasizes these shared classes while diminishing the significance of distinctive features [6].
The MulSupCon method employs a positive sample identification mechanism analogous to ANY, samples $p_{j}$, where $j\in{3,4,5}$ are designated as positive instances. However, MulSupCon distinguishes itself by evaluating each label individually and forming multiple positive sets for a single anchor sample. This approach aggregates positive samples based on the number of overlapping labels between the positive samples and the anchor, thereby expanding the space of positive sets:
$\mathcal{P}(i)=\{p_{2},p_{2},p_{2},p_{3},p_{4},p_{4},p_{5},p_{5},p_{5}\}$
Subsequently, the loss for $p_{j}$ in Eq. (6) are as follows by Lemma 1:
$\mathcal{L}(R2)\approx\mathcal{L}(R5)\neq\mathcal{L}(R3)\neq\mathcal{L}(R4)$
For this example (see Figure 1), the MulSupCon successfully discriminates $R3$ and $R4$ from $R2$ and $R5$; however, it fails to establish a distinction between $R2$ and $R5$. This limitation arises primarily because MulSupCon exclusively considers the overlapping regions (Similarity  [^3] The definition of Similarity is introduced in Section II-D) between anchor $i$ and sample $p$ (i.e., The intersection of sets $\mathcal{S}$ and $\mathcal{T}$), while disregarding the complementary non-intersecting domains (Dissimilarity  [^4] The definition of Dissimilarity is introduced in Section II-D). That is to say, the similarity between positive samples and anchors is considers, but not yet dissimilarity, which is one of critical information for representation learning in MSCL.
Leveraging the proposed multi-label relations, our theoretical analysis systematically elucidates the limitations of existing methods and establishes a rigorous foundation for investigating the profound exploration of concepts of similarity and dissimilarity, and the design of contrastive loss function.
II-D Similarity-Dissimilarity Loss
To address the aforementioned challenges, we introduce the concepts of similarity and dissimilarity based on set-theoretic relations: (i) As depicted in Figure 1, Similarity represents the intersection of sets (i.e., $\mathcal{S}\cap\mathcal{T}$), and (ii) we define Dissimilarity as the set difference between $\mathcal{T}$ and the intersection $\mathcal{S}\cap\mathcal{T}$ with respect to sample $p$ (i.e., $\mathcal{T}-\mathcal{S}\cap\mathcal{T}$). For each anchor $i$, we formulate the Similarity-Dissimilarity Loss as:
$\mathcal{L}_{i}^{our}=\frac{-1}{|\mathcal{P}(i)|}\sum_{p\in\mathcal{P}(i)}\log\frac{\mathcal{K}_{i,p}^{s}\mathcal{K}_{i,p}^{d}\exp(\boldsymbol{z}_{i}\cdot\boldsymbol{z}_{p}/\tau)}{\sum_{a\in\mathcal{A}(i)}\exp(\boldsymbol{z}_{i}\cdot\boldsymbol{z}_{a}/\tau)}$ (13)
Here, we define $\mathcal{K}_{i,p}^{s}$ and $\mathcal{K}_{i,p}^{d}$ that quantify the Similarity and Dissimilarity factors for a given anchor $i$ and a positive sample $p$, respectively. These factors are formally defined as follows:
$\mathcal{K}_{i,p}^{s}=\frac{|\tilde{\boldsymbol{y}}_{p}^{s}|}{|\tilde{\boldsymbol{y}}_{i}|}=\frac{|\mathcal{S}\cap\mathcal{T}|}{|\mathcal{S}|}$ (14)
and
$\mathcal{K}_{i,p}^{d}=\frac{1}{1+|\tilde{\boldsymbol{y}}_{p}^{d}|}=\frac{1}{1+|\mathcal{T}\setminus(\mathcal{S}\cap\mathcal{T})|}$ (15)
where we define the following set-theoretic quantities:
  • $|\tilde{\boldsymbol{y}}_{i}|=|\mathcal{S}|$ denotes the cardinality of the label space $\tilde{\boldsymbol{y}}_{i}$.
  • $|\tilde{\boldsymbol{y}}_{p}^{s}|=|\mathcal{S}\cap\mathcal{T}|$ measures the cardinality of the intersection of sets $\mathcal{S}$ and $\mathcal{T}$.
  • $|\tilde{\boldsymbol{y}}_{p}^{d}|=|\mathcal{T}\setminus(\mathcal{S}\cap\mathcal{T})|$ represents the cardinality of the relative complement with respect to sample $p$.
The product of $\mathcal{K}_{i,p}^{s}$ and $\mathcal{K}_{i,p}^{d}$ is termed as similarity-dissimilarity factor. Moreover, the following relation holds:
$|\tilde{\boldsymbol{y}}_{p}^{d}|=|\tilde{\boldsymbol{y}}_{p}|-|\tilde{\boldsymbol{y}}_{p}^{s}|\geq 0$ (16)
where $|\tilde{\boldsymbol{y}}_{p}|$ represents the cardinality of the label space associated with sample $p$.
Specifically, the Similarity-Dissimilarity Loss Loss reduces to Eq. (3), when the following conditions are simultaneously satisfied:
[…]
Accordingly, our proposed loss function constitutes a generalized form of the basic supervised contrastive loss (see Eq. (3)). In particular, Eq. (3) represents a particular case of the Similarity-Dissimilarity Loss. Moreover, our contrastive loss unifies both single-label and multi-label supervised contrastive loss functions within a comprehensive form and paradigm.
II-E Case Analysis
Let us examine the behavior of our loss function through a detailed analysis of five distinct relational cases illustrated in Figure 1. Consider the following sequences of cardinalities:
[…]
Applying these values to Eq. (14) and (15), we obtain:
[…]
Consequently, the product of these measures yields:
$\mathcal{K}_{i,p}^{s}\mathcal{K}_{i,p}^{d}=\{0,1,\frac{1}{9},\frac{2}{3},\frac{1}{3}\}$
When evaluating Eq. (13), these distinct relations ($R2$ through $R5$) generate unique loss values, establishing the following inequalities:
$\mathcal{L}(R2)\neq\mathcal{L}(R3)\neq\mathcal{L}(R4)\neq\mathcal{L}(R5)$
The proposed loss function effectively discriminates among the five distinct relations through a principled re-weighting mechanism, as formulated in Eq. (13), (14), and (15), comparing to existing methods in MSCL.
Furthermore, in contrast to MulSupCon, the Similarity-Dissimilarity Loss preserves the cardinality of positive sets while maintaining computational efficiency, as it requires no additional computational overhead.
II-F Theoretical Analysis
The proposed loss function incorporates a weighting mechanism through the product of factors $\mathcal{K}_{i,p}^{s}$ and $\mathcal{K}_{i,p}^{d}$. By construction, the similarity-dissimilarity factor $\mathcal{K}_{i,p}^{s}\mathcal{K}_{i,p}^{d}$ is constrained to the closed interval $[0,1]$ across all possible relational configurations. Hence, it is written as:
$\mathcal{K}_{i,p}^{s}\mathcal{K}_{i,p}^{d}\in[0,1]$ (18)
For notational conciseness, let us denote the product of Similarity and Dissimilarity factors across the five relations as $\{\mathcal{K}_{m}^{s}\mathcal{K}_{m}^{d}\}_{m=1,2,3,4,5}$.
Theorem 1.
Let $\mathcal{K}_{m}^{s}$ and $\mathcal{K}_{m}^{d}$ be the Similarity and Dissimilarity operators, respectively, as defined in Eq. (14) and (15). For the case $m=1$, their product vanishes:
$\mathcal{K}_{m}^{s}\mathcal{K}_{m}^{d}=0,\quad\text{when }m=1$ (19)
Proof.
Consider the case where $m=1$. By definition, we have $\mathcal{S}\cap\mathcal{T}=\varnothing$. This implies:
$|\tilde{\boldsymbol{y}}_{p}^{s}|$ $=|\mathcal{S}\cap\mathcal{T}|=|\varnothing|=0$
$\therefore\mathcal{K}_{1}^{s}$ $=\frac{|\tilde{\boldsymbol{y}}_{p}^{s}|}{|\tilde{\boldsymbol{y}}_{i}|}=\frac{0}{|\tilde{\boldsymbol{y}}_{i}|}=0$
Since $\mathcal{K}_{1}^{s}=0$ and $\mathcal{K}_{1}^{d}$ is finite by construction, we conclude:
$\mathcal{K}_{1}^{s}\mathcal{K}_{1}^{d}=0\cdot\mathcal{K}_{1}^{d}=0\quad$ (20)
Theorem 2.
Consider the Similarity operator $\mathcal{K}_{m}^{s}$ and Dissimilarity operator $\mathcal{K}_{m}^{d}$ as defined in Eq. (14) and (15). For the case $m=2$, their product equals unity:
$\mathcal{K}_{m}^{s}\mathcal{K}_{m}^{d}=1,\quad\text{when }m=2$ (21)
Proof.
Consider the case where $m=2$. By hypothesis, we have $\mathcal{S}=\mathcal{T}$. This equality implies:
$\mathcal{K}_{2}^{s}$ $=\frac{|\mathcal{S}\cap\mathcal{T}|}{|\mathcal{S}|}=\frac{|\mathcal{S}|}{|\mathcal{S}|}=1$
$\mathcal{K}_{2}^{d}$ $=\frac{1}{1+|\mathcal{T}\setminus(\mathcal{S}\cap\mathcal{T})|}=\frac{1}{1+|\varnothing|}=1$
where we have used the fact that $\mathcal{T}\setminus(\mathcal{S}\cap\mathcal{T})=\varnothing$ when $\mathcal{S}=\mathcal{T}$. Thus, we conclude:
$\mathcal{K}_{2}^{s}\mathcal{K}_{2}^{d}=1\cdot 1=1\quad$ (22)
Theorem 3.
Let $\mathcal{K}_{m}^{s}$ and $\mathcal{K}_{m}^{d}$ be the Similarity and Dissimilarity operators as defined in Eq. (14) and (15), respectively. For $m\in\{3,4,5\}$, their product is strictly bounded between 0 and 1:
$10<\mathcal{K}_{m}^{s}\mathcal{K}_{m}^{d}<$ 1 (23)
Proof.
Consider $m\in\{3,4,5\}$. Under these cases, we have:
[…]
$\mathcal{S}$ $\neq\mathcal{T}$ (25)
We first establish the strict positivity. Given $|\mathcal{S}|>0$ and conditions (24)-(25), we have:
$\mathcal{K}_{m}^{s}$ $=\frac{|\mathcal{S}\cap\mathcal{T}|}{|\mathcal{S}|}>0$
$\mathcal{K}_{m}^{d}$ $=\frac{1}{1+|\mathcal{T}\setminus(\mathcal{S}\cap\mathcal{T})|}>0$
For the upper bound, we consider three cases:
Case 1 ($m=3$): By Eq. (10), we have three conditions: $\mathcal{S}\cap\mathcal{T}\neq\varnothing$, $\mathcal{S}\nsubseteq\mathcal{T}$, and $\mathcal{T}\nsubseteq\mathcal{S}$. These conditions lead to:
$|\mathcal{S}\cap\mathcal{T}|$ $<|\mathcal{S}|\implies\mathcal{K}_{3}^{s}<1$
$|\mathcal{T}\setminus(\mathcal{S}\cap\mathcal{T})|$ $>0\implies\mathcal{K}_{3}^{d}<1$
Therefore, $\mathcal{K}_{3}^{s}\mathcal{K}_{3}^{d}<1$.
Case 2 ($m=4$): When $m=4$, by Eq. (11), we have $\mathcal{S}\supseteq\mathcal{T}$. This subset relation implies:
$\mathcal{K}_{4}^{s}$ $=\frac{|\mathcal{S}\cap\mathcal{T}|}{|\mathcal{S}|}=\frac{|\mathcal{T}|}{|\mathcal{S}|}<1$
$\mathcal{K}_{4}^{d}$ $=\frac{1}{1+|\mathcal{T}\setminus(\mathcal{S}\cap\mathcal{T})|}=\frac{1}{1+|\varnothing|}=1$
where the strict inequality $\mathcal{K}_{4}^{s}<1$ follows from $|\mathcal{T}|<|\mathcal{S}|$ (since $\mathcal{S}\supsetneq\mathcal{T}$), and $\mathcal{K}_{4}^{d}=1$ is a consequence of $\mathcal{T}\setminus(\mathcal{S}\cap\mathcal{T})=\varnothing$ when $\mathcal{S}\supsetneq\mathcal{T}$. Therefore:
$\mathcal{K}_{4}^{s}\mathcal{K}_{4}^{d}=\mathcal{K}_{4}^{s}\cdot 1=\mathcal{K}_{4}^{s}<1$
Case 3 ($m=5$): When $m=5$, by Eq. (12), we have $\mathcal{S}\subsetneq\mathcal{T}$. This subset relation implies:
$\mathcal{K}_{5}^{s}$ $=\frac{|\mathcal{S}\cap\mathcal{T}|}{|\mathcal{S}|}=\frac{|\mathcal{S}|}{|\mathcal{S}|}=1$
$\mathcal{K}_{5}^{d}$ $=\frac{1}{1+|\mathcal{T}\setminus(\mathcal{S}\cap\mathcal{T})|}=\frac{1}{1+|\mathcal{T}\setminus\mathcal{S}|}<1$
where $\mathcal{K}_{5}^{s}=1$ follows from the fact that $\mathcal{S}\cap\mathcal{T}=\mathcal{S}$ when $\mathcal{S}\subsetneq\mathcal{T}$. The strict inequality $\mathcal{K}_{5}^{d}<1$ holds because:
$\mathcal{S}\subsetneq\mathcal{T}$ $\implies|\mathcal{T}\setminus\mathcal{S}|>0$
$\implies 1+|\mathcal{T}\setminus\mathcal{S}|>1$
$\implies\frac{1}{1+|\mathcal{T}\setminus\mathcal{S}|}<1$
Therefore, we can conclude:
$\mathcal{K}_{5}^{s}\mathcal{K}_{5}^{d}=1\cdot\mathcal{K}_{5}^{d}=\mathcal{K}_{5}^{d}<1$
Combining the results with Propositions 1 and 2, we obtain complete ordering for all $m\in\{1,2,3,4,5\}$. The products $\mathcal{K}_{m}^{s}\mathcal{K}_{m}^{d}$ satisfy:
$0=\mathcal{K}_{1}^{s}\mathcal{K}_{1}^{d}<\mathcal{K}_{m}^{s}\mathcal{K}_{m}^{d}<\mathcal{K}_{2}^{s}\mathcal{K}_{2}^{d}=1,\quad m\in\{3,4,5\}\quad$ (26)
Based on Theorem 1, 2, and 3, the product of weighting factors $\mathcal{K}_{i,p}^{s}$ and $\mathcal{K}_{i,p}^{d}$ is bounded within the interval $[0,1]$, which aligns with fundamental principles of loss functions and set-theoretic relations. The non-negative lower bound adheres to the essential property of loss functions being strictly positive [9]. Given that our proposed loss function generalizes the supervised contrastive loss [5] and incorporates multi-label relation definitions, the upper bound naturally equals $1$. Furthermore, this mathematical framework demonstrates that our proposed contrastive loss can dynamically adjust the weighting factor within $[0,1]$, effectively differentiating sample features with rigorous mathematical justification for both the formulation and efficacy of the loss function.
Theorem 4.
Let $i\in\mathcal{I}$ be a fixed anchor sample, and let $p_{3},p_{4}\in\mathcal{P}(i)$ be positive samples corresponding to relations $R_{3}$ and $R_{4}$, respectively. Suppose their label spaces satisfy the cardinality constraint:
$|\tilde{\boldsymbol{y}}_{p_{3}}|=|\tilde{\boldsymbol{y}}_{p_{4}}|$ (27)
Then, the product of similarity and dissimilarity operators satisfies the strict inequality:
$\mathcal{K}_{4}^{s}\mathcal{K}_{4}^{d}>\mathcal{K}_{3}^{s}\mathcal{K}_{3}^{d}$ (28)
Proof.
Let us establish the strict inequality $\mathcal{K}_{4}^{s}\mathcal{K}_{4}^{d}>\mathcal{K}_{3}^{s}\mathcal{K}_{3}^{d}$ through direct comparison. From definitions (14) and (15), we have:
$\mathcal{K}_{4}^{s}\mathcal{K}_{4}^{d}=\frac{|\tilde{\boldsymbol{y}}_{p_{4}}|}{|\tilde{\boldsymbol{y}}_{i}|}>\mathcal{K}_{3}^{s}\mathcal{K}_{3}^{d}=\frac{|\tilde{\boldsymbol{y}}_{p_{3}}-\tilde{\boldsymbol{y}}_{p_{3}}^{d}|}{|\tilde{\boldsymbol{y}}_{i}|}\cdot\frac{1}{1+|\tilde{\boldsymbol{y}}_{p_{3}}^{d}|}$
$\Rightarrow$
$\frac{|\tilde{\boldsymbol{y}}_{p_{4}}|(1+|\tilde{\boldsymbol{y}}_{p_{3}}^{d}|)}{|\tilde{\boldsymbol{y}}_{i}|(1+|\tilde{\boldsymbol{y}}_{p_{3}}^{d}|)}>\frac{|\tilde{\boldsymbol{y}}_{p_{4}}-\tilde{\boldsymbol{y}}_{p_{3}}^{d}|}{|\tilde{\boldsymbol{y}}_{i}|(1+|\tilde{\boldsymbol{y}}_{p_{3}}^{d}|)}$
$\Rightarrow$
$|\tilde{\boldsymbol{y}}_{p_{4}}|(1+|\tilde{\boldsymbol{y}}_{p_{3}}^{d}|)>|\tilde{\boldsymbol{y}}_{p_{3}}-\tilde{\boldsymbol{y}}_{p_{3}}^{d}|$
By the cardinality constraint (27) in the theorem:
$|\tilde{\boldsymbol{y}}_{p_{3}}|(1+|\tilde{\boldsymbol{y}}_{p_{3}}^{d}|)>|\tilde{\boldsymbol{y}}_{p_{3}}-\tilde{\boldsymbol{y}}_{p_{3}}^{d}|$
where the strict inequality follows from the fact that for any positive real numbers $a,b>0$:
$a(1+b)>a-b$
This inequality holds trivially, thereby establishing the original claim $\mathcal{K}_{4}^{s}\mathcal{K}_{4}^{d}>\mathcal{K}_{3}^{s}\mathcal{K}_{3}^{d}$. ∎
Theorem 5.
Let $i\in\mathcal{I}$ be a fixed anchor sample, and let $p_{3},p_{5}\in\mathcal{P}(i)$ be positive samples corresponding to relations $R_{3}$ and $R_{5}$, respectively. Suppose:
$|\tilde{\boldsymbol{y}}_{p_{5}}^{d}|\leq|\tilde{\boldsymbol{y}}_{p_{3}}^{d}|$ (29)
Then, the product of Similarity and Dissimilarity operators satisfies the strict inequality:
$\mathcal{K}_{5}^{s}\mathcal{K}_{5}^{d}>\mathcal{K}_{3}^{s}\mathcal{K}_{3}^{d}$ (30)
Proof.
From definitions (14) and (15), we have:
$\mathcal{K}_{3}^{s}\mathcal{K}_{3}^{d}$ $=\frac{|\tilde{\boldsymbol{y}}_{p_{3}}-\tilde{\boldsymbol{y}}_{p_{3}}^{d}|}{|\tilde{\boldsymbol{y}}_{i}|}\cdot\frac{1}{1+|\tilde{\boldsymbol{y}}_{p_{3}}^{d}|}$
$\mathcal{K}_{5}^{s}\mathcal{K}_{5}^{d}$ $=\frac{1}{1+|\tilde{\boldsymbol{y}}_{p_{5}}^{d}|}$
Taking the ratio:
$\frac{\mathcal{K}_{5}^{s}\mathcal{K}_{5}^{d}}{\mathcal{K}_{3}^{s}\mathcal{K}_{3}^{d}}=\frac{|\tilde{\boldsymbol{y}}_{i}|(1+|\tilde{\boldsymbol{y}}_{p_{3}}^{d}|)}{|\tilde{\boldsymbol{y}}_{p_{3}}-\tilde{\boldsymbol{y}}_{p_{3}}^{d}|(1+|\tilde{\boldsymbol{y}}_{p_{5}}^{d}|)}$
By the properties of cardinality and set difference:
$|\tilde{\boldsymbol{y}}_{p_{3}}-\tilde{\boldsymbol{y}}_{p_{3}}^{d}|\leq|\tilde{\boldsymbol{y}}_{i}|$
Given the constraint (29), $|\tilde{\boldsymbol{y}}_{p_{5}}^{d}|\leq|\tilde{\boldsymbol{y}}_{p_{3}}^{d}|$, we have:
$\frac{|\tilde{\boldsymbol{y}}_{i}|(1+|\tilde{\boldsymbol{y}}_{p_{3}}^{d}|)}{|\tilde{\boldsymbol{y}}_{p_{3}}-\tilde{\boldsymbol{y}}_{p_{3}}^{d}|(1+|\tilde{\boldsymbol{y}}_{p_{5}}^{d}|)}>1$
Therefore, $\mathcal{K}_{5}^{s}\mathcal{K}_{5}^{d}>\mathcal{K}_{3}^{s}\mathcal{K}_{3}^{d}$. ∎
Theorem 4 and 5 establish strict dominance relations between relation types $R3$, $R4$, and $R5$, demonstrating that $\mathcal{K}_{4}^{s}\mathcal{K}_{4}^{d}>\mathcal{K}_{3}^{s}\mathcal{K}_{3}^{d}$ when $|\tilde{\boldsymbol{y}}_{p_{3}}|=|\tilde{\boldsymbol{y}}_{p_{4}}|$ and $\mathcal{K}_{5}^{s}\mathcal{K}_{5}^{d}>\mathcal{K}_{3}^{s}\mathcal{K}_{3}^{d}$ when $|\tilde{\boldsymbol{y}}_{p_{5}}^{d}|\leq|\tilde{\boldsymbol{y}}_{p_{3}}^{d}|$. These inequalities, proved through careful mathematical derivation using set cardinality properties and fundamental principles of real analysis, reveal a well-defined hierarchical structure in the weighting factors. This hierarchical relations ensures that our loss function appropriately modulates the contribution of different relation types during the learning process, providing theoretical guarantees for the effectiveness of our proposed approach in capturing complex relations within the data.
Our theoretical analysis establishes a comprehensive mathematical foundation for the proposed loss function through five key theorems. These theoretical guarantees, derived through rigorous set-theoretic analysis, demonstrate that our loss function effectively modulates the contribution of different relation types while maintaining proper mathematical bounds, thereby providing a solid theoretical foundation for its application in multi-label contrastive learning.
TABLE I: Statistics of datasets.
Dataset Train Val Test Total # labels Avg # labels
MS-COCO 82.0k 20.2k 20.2k 80 2.9
PASCAL 5.0k 2.5k 2.5k 20 1.5
NUS-WIDE 125.4k 41.9k 41.9k 81 2.4
AAPD 37.8k 6.7k 11.3k 54 2.4
MIMIC-III-Full 47,723 1,631 3,372 8,692 15.7
MIMIC-III-50 8,066 1,573 1,729 50 5.7
MIMIC-IV-ICD9-Full 188,533 7,110 13,709 11,145 13.4
MIMIC-IV-ICD9-50 170,664 6,406 12,405 50 4.7
MIMIC-IV-ICD10-Full 110,442 4,017 7,851 25,230 16.1
MIMIC-IV-ICD10-50 104,077 3,805 7,368 50 5.4
III Experiments
The previous theoretical analysis establishes a rigorous mathematical foundation for our method, validating both the formulation and efficacy of the proposed loss function. In our experimental evaluation, we focus on assessing the effectiveness and robustness of Similarity-Dissimilarity Loss in the MSCL framework. Rather than comparing with other multi-label classification approaches, we emphasize that Similarity-Dissimilarity Loss primarily aims to enable models to learn generalizable and transferable features that enhance performance across diverse downstream tasks (classification, detection, and clustering) instead of optimizing for any specific task. We conduct the experiments to compare Similarity-Dissimilarity Loss with current contrastive loss functions (ALL, ANY, and MulSupCon) in a comprehensive evaluation, considering: (i) Data modality: image and text data; (ii) Domain-specific: general text data (AAPD) and medical domain (MIMIC III and IV); (iii) Data distribution: full setting (extreme long-tailed distribution) and top-50 frequent labels setting; (iv) ICD code versions: ICD-9 and ICD-10, and (v) Models: ResNet-50, RoBERTa-based, Llama-3.1-8B, and PLM-ICD.
III-A Datasets and Metrics
To rigorously evaluate the efficacy of our proposed loss function, we conducted comprehensive experiments across three distinct data modalities: visual data, textual data, and specialized medical corpus data (MIMIC datasets). The MIMIC datasets are particularly noteworthy for their exceptionally large label space and pronounced long-tailed distributions [10]. This long-tailed characteristic, which is especially prevalent in multi-label classification scenarios, facilitates a robust assessment of the performance of our loss function across heterogeneous data distributions. Comprehensive statistical analyses of all experimental datasets are presented in Table I.
  • MS-COCO (Microsoft Common Objects in Context) [11] consists of over 330,000 images annotated across 80 object categories, providing rich semantic information for object detection, segmentation, and captioning tasks that has significantly advanced computer vision research since its introduction by Microsoft.
  • PASCAL VOC [12] contains 9,963 natural images with standardized annotations spanning 20 object categories, enabling rigorous evaluation of classification, detection, and segmentation algorithms in computer vision.
  • NUS-WIDE [13] is a large-scale web image collection comprising approximately 269,000 Flickr images annotated with 81 concept categories and user tags, widely used as a benchmark for multi-label image classification.
  • AAPD (Arxiv Academic Paper Dataset) [14] is a text corpus containing 55,840 scientific paper abstracts from arXiv with multi-label annotations across various subject categories, designed specifically for benchmarking multi-label text classification and document categorization algorithms.
  • MIMIC-III  [^5] We are granted access to MIMIC-III Clinical Database (v1.4) [15] includes records labeled with expert-annotated ICD-9 codes, which identify diagnoses and procedures. We adhere to the same splits as in previous works [16], employing two settings: MIMIC-III-Full, which includes all ICD-9 codes, and MIMIC-III-50, which includes only the 50 most frequent codes.
  • MIMIC-IV  [^6] We are granted access to MIMIC-IV (v2.2) [17] contains records annotated with both ICD-9 and ICD-10 codes, where each code is subdivided into sub-codes that often capture specific circumstantial details. we follow prior studies [18] and utilize four settings: MIMIC-IV-ICD9-Full, MIMIC-IV-ICD9-50, MIMIC-IV-ICD10-Full, and MIMIC-IV-ICD10-50.
Metrics. Consistent with prior research [16, 18], we report macro/micro-AUC, macro/micro-F1, and precision at K (P@$\mathcal{K}$) metrics on MIMIC datasets, where $\mathcal{K}=\{5,8\}$ for different settings. Moreover, micro/macro-F1 and mAP are used for image datasets following [8, 6, 19].
III-B Baseline Loss Functions and Encoders
This study evaluates the proposed Similarity-Dissimilarity Loss in comparison with three established baseline loss functions: (i) ALL, (ii) ANY, and (iii) MulSupCon [6], all implemented within the MSCL framework.
For experimental evaluation, we employ modality-specific encoder architectures tailored to each data type. For image data, ResNet-50 [20] serves as the encoder architecture, consistent with established methodologies [8, 7, 6]. For textual data, we utilize pre-trained large language models (LLMs), specifically RoBERTa-base [21] and Llama-3.1-8B [22] with Low-Rank Adaptation (LoRA) [23]. Additionally, for the specialized task of ICD coding on MIMIC datasets, we implement PLM-ICD [24], a model specifically designed for ICD coding using LLMs.
III-C Implementation Details
Within the MSCL framework, we implement a two-phase training method as established by Khosla [5]: (i) encoder training, wherein the model learns to generate vector representations that maximize similarity between instances of the same class while distinguishing them from other classes; and (ii) classifier training, which utilizes the trained encoder and freeze it to train the classifier.
In the representation training, we use a standard cosine learning rate scheduler with a $0.05$ warm-up period and set the temperature $\tau=0.07$. The projection head comprises two MLP layers with ReLU activation function and employs contrastive loss function for the training, where the projected representation $\boldsymbol{z}_{k}=Proj(Enc(\tilde{\boldsymbol{x}}_{k}))\in\mathbb{R}^{D_{P}}$. Here $h=Enc(\tilde{\boldsymbol{x}}_{k})$ denotes the encoded feature vectors and the projection dimension $D_{P}=256$. For subsequent classifier training, the projection head is removed, a linear layer is appended to the frozen encoder, and binary cross-entropy (BCE) loss is utilized for optimization.
For image data, we employ ResNet-50 using stochastic gradient descent (SGD) with momentum. The input images are set up at a resolution of $224\times 224$ pixels. For text data, RoBERTa-base and Llama-3.1-8B serve as backbone encoders implemented via Huggingface platform [25]. RoBERTa configures with a dropout rate of $0.1$ and AdamW optimizer with a weight decay of $0.01$, exempting bias and LayerNorm from weight decay. Compared with full-parameter fine-tuning, we employ LoRA [23] to efficiently fine-tune large model Llama. LoRA configures with the low-rank dimension $r=16$, scaling factor $\alpha=32$ and dropout as $0.1$. There is no KV cache to save memory during training. To enhance computational efficiency, BFloat16 precision is used for the training. The hyperparameters and detailed configuration are shown our code  [^7] https://github.com/guangminghuang/similarity-dissimilarity-loss.
TABLE II: Results on image datasets. We compare our method with baselines (ALL, ANY and MulSupCon) on MS-COCO, PASCAL and NUS-WIDE. mAP standards for mean Average Precision.
Method MS-COCO PASCAL NUS-WIDE
Micro-F1 Macro-F1 mAP Micro-F1 Macro-F1 mAP Micro-F1 Macro-F1 mAP
ALL 68.93 63.32 64.11 82.53 79.87 79.32 70.25 52.84 51.35
ANY 64.80 57.37 56.90 82.31 79.65 79.15 68.42 50.65 49.28
MulSupCon 71.33 66.25 67.69 82.75 80.26 79.58 71.88 54.36 52.47
Ours 73.40 70.03 69.20 83.63 81.10 79.75 73.35 57.49 56.74
Figure 2: Comparison of performance improvements between Similarity-Dissimilarity Loss and MulSupCon.
Figure 3: Comparison standard deviation of image datasets on micro-F1, macro-F1 and mAP metrics.
IV Results and Analysis
IV-A Evaluation on Image
The experimental results in Table II demonstrate that our proposed loss function outperforms baselines across all metrics, including micro-F1, macro-F1, and mAP, on all image datasets (MS-COCO, PASCAL, and NUS-WIDE). Compared to MulSulCon, Similarity-Dissimilarity Loss achieves significant improvements of $2.07/3.78/1.51$ in Micro-F1, Macro-F1, and mAP on MS-COCO and $1.47/3.13/4.27$ on NUS-WIDE.
Figure 2 illustrates the comparison between Similarity-Dissimilarity Loss and MulSupCon as measured by micro- and macro-F1 metrics. The results indicate that our method yields substantially greater improvements in macro-F1 compared to micro-F1 across all image datasets. Specifically, macro-F1 increases by $5.7\%$ on MS-COCO and $5.8\%$ on NUS-WIDE, whereas micro-F1 exhibits more modest improvements of $2.9\%$ and $2.0\%$, respectively. Macro-F1 assigns equal importance to each class regardless of its frequency, rendering it particularly appropriate for evaluating performance on imbalanced datasets where minority class prediction accuracy is critical [9, 3]. In contrast, micro-F1 places more considerable weight on classes with more samples, making it more appropriate when larger classes should have a more potent influence on the overall score [9, 26]. Multi-label classification inherently faces more pronounced challenges with long-tailed distributions than single-label classification due to exponential output space complexity, intricate label co-occurrence patterns, and high annotation costs [3]. The observed superior improvement in macro-F1 metrics provides compelling evidence that our method demonstrates exceptional efficacy in addressing long-tailed distribution challenges, a capability particularly crucial in multi-label scenarios.
However, on the PASCAL dataset, our method demonstrates mere marginal improvements, with gains of $0.88/0.84/0.17$ in micro/macro-F1/mAP, respectively. This limited enhancement can be attributed to the structural characteristics of PASCAL, wherein the average number of labels per instance is approximately $1.5$ (as detailed in Table I), causing the task to approximate single-label classification, particularly when the batch size is limited [5]. Consequently, loss functions specifically designed for multi-label scenarios exert minimal influence on model performance under these conditions. As Audibert et al. and [19] have demonstrated, the cardinality of the label space constitutes a significant determinant of model efficacy within MSCL .
Furthermore, Figure 3 reveals that the standard deviation across four methods for PASCAL equals $0.58/0.64/0.27$ in micro/macro-F1/mAP, which are considerably lower than the corresponding standard deviations observed for the MS-COCO and NUS-WIDE. This statistical finding suggests that the efficacy of specialized multi-label loss functions diminishes significantly when the average label cardinality per instance approaches $1$ in MSCL. This finding further corroborates our theoretical analysis and hypothesis in the Section II, wherein Similarity-Dissimilarity Loss degenerates to single-label scenarios (see Eq. (17)).
TABLE III: Results on MIMIC-III-Full, MIMIC-IV-ICD9-Full and MIMIC-IV-ICD10-Full test sets. The best scores among backbone encoder models are marked in bold.
MIMIC-III-Full MIMIC-IV-ICD9-Full MIMIC-IV-ICD10-Full
AUC F1 AUC F1 AUC F1
Method Macro Micro Macro Micro P@8 Macro Micro Macro Micro P@8 Macro Micro Macro Micro P@8
RoBERTa
ALL 89.87 95.83 7.94 53.08 71.06 93.04 98.57 11.76 57.73 64.75 89.46 98.19 4.23 53.82 64.17
ANY 88.15 94.18 7.13 51.35 68.92 92.86 98.14 11.17 57.42 64.48 89.09 98.07 4.02 52.28 62.65
MulSupCon 90.37 96.38 8.64 54.16 71.24 93.87 99.34 12.83 58.67 65.89 90.53 98.74 4.56 54.09 65.46
Ours 90.78 96.67 9.19 54.63 71.38 94.13 99.36 13.08 58.85 66.29 90.68 98.86 4.72 54.89 66.07
Llama
ALL 91.27 96.94 8.38 54.75 72.63 94.52 98.93 12.34 58.97 66.35 90.78 98.57 4.53 54.98 65.31
ANY 90.64 96.38 7.82 53.97 71.85 94.19 98.74 11.93 58.68 65.92 90.36 98.32 4.37 54.24 64.78
MulSupCon 91.68 97.23 8.79 55.36 72.94 94.87 99.42 12.96 59.35 66.73 91.15 98.97 4.72 55.29 66.16
Ours 91.93 97.57 9.26 55.87 73.28 95.14 99.58 13.37 59.69 67.14 91.38 99.15 4.94 55.67 66.59
PLM-ICD
ALL 92.58 98.69 10.73 60.06 76.84 96.95 99.28 14.18 62.83 70.53 91.87 98.79 4.83 57.36 69.29
ANY 91.09 97.36 9.24 58.87 75.38 95.85 98.17 12.64 61.82 69.58 90.54 97.72 4.54 55.86 68.17
MulSupCon 93.46 99.13 11.68 61.42 77.65 97.86 99.32 14.47 64.23 71.97 92.83 99.38 5.43 58.15 70.19
Ours 94.47 99.43 12.46 62.34 78.42 98.47 99.59 15.04 64.95 72.95 93.75 99.57 5.74 58.76 70.79
IV-B Evaluation on Text
TABLE IV: Results on AAPD Dataset. We compare our proposed Similarity-Dissimilarity Loss with baselines on general text data using RoBERTa-based and Llama-3.1-8B models
Method RoBERTa Llama
Micro-F1 Macro-F1 Micro-F1 Macro-F1
ALL 73.23 59.41 74.32 60.47
ANY 72.31 58.55 73.41 59.63
MulSupCon 73.64 60.52 74.72 61.58
Ours 74.54 62.31 75.61 63.42
We further evaluate our method on general text data, and the results demonstrate that our proposed loss function consistently surpasses baseline methods for both RoBERTa and Llama models across all metrics on the AAPD dataset (See Table IV). In contrast to the significant performance gains observed on image data, Similarity-Dissimilarity Loss achieves more modest enhancements of $0.90/1.79$ in micro/macro-F1 scores on RoBERTa, and $0.89/1.84$ on Llama. This attenuated performance differential can be attributed to the extensive knowledge already encoded within LLMs through their comprehensive pre-training paradigms [27].
Moreover, as illustrated in Figure 4, performance variations of contrastive loss functions for MSCL on both RoBERTa and Llama models are relatively minimal. Specifically, the standard deviations in micro-F1 are $0.80$ and $0.79$ on RoBERTa and Llama, respectively, while the corresponding standard deviations for macro-F1 metrics are $1.41$ and $1.42$. Unlike image classification in MSCL paradigm, performance improvements in text classification are predominantly attributable to the intrinsic representational capabilities of model architecture of LLMs. Consequently, while fine-tuning the pre-trained weights of LLMs during the contrastive learning phase can yield marginal performance improvements, this methodological approach demonstrates substantially greater efficacy for visual classification tasks compared to textual classification.
Figure 4: Comparison of RoBERTa and Llama across micro- and macro-F1 on AAPD Dataset.
IV-C Evaluation on Medical Domain
TABLE V: Results on MIMIC-III-50, MIMIC-IV-ICD9-50 and MIMIC-IV-ICD10-50 test sets. The best scores among backbone encoder models are marked in bold.
MIMIC-III-50 MIMIC-IV-ICD9-50 MIMIC-IV-ICD10-50
AUC F1 AUC F1 AUC F1
Method Macro Micro Macro Micro P@5 Macro Micro Macro Micro P@5 Macro Micro Macro Micro P@5
RoBERTa
ALL 87.73 90.57 57.38 61.84 61.29 93.84 94.46 67.63 72.24 60.92 91.43 93.52 64.86 67.65 60.07
ANY 87.36 89.42 56.25 60.83 60.32 93.37 93.73 67.39 71.97 60.24 90.06 92.03 64.09 66.54 58.08
MulSupCon 88.02 91.24 57.83 62.26 61.53 94.73 95.28 68.63 73.32 61.98 92.09 93.95 65.43 68.54 61.36
Ours 88.86 93.14 60.03 62.43 62.06 94.92 95.43 69.05 73.54 62.23 92.43 94.34 66.07 70.24 62.09
Llama
ALL 88.93 91.67 60.32 64.58 62.87 94.32 95.28 69.18 73.72 61.42 92.35 94.57 66.38 69.83 61.75
ANY 88.57 91.09 59.72 64.03 62.19 94.05 94.85 68.79 73.19 61.07 91.89 93.97 65.82 69.09 61.12
MulSupCon 89.21 92.13 60.85 65.12 63.23 94.74 95.83 69.76 74.46 61.95 92.73 95.12 66.85 70.57 62.43
Ours 89.54 92.49 61.32 65.67 63.69 94.97 96.07 70.21 74.87 62.32 93.07 95.53 67.23 71.23 62.81
PLM-ICD
ALL 90.13 93.02 65.18 69.43 65.26 95.18 96.42 71.31 75.83 62.45 93.53 95.97 68.96 73.14 64.52
ANY 89.03 92.07 63.73 68.14 63.84 93.73 95.34 70.23 74.43 61.42 92.27 94.42 67.95 71.83 63.17
MulSupCon 91.23 94.04 66.17 70.32 66.42 96.32 97.63 72.64 76.93 63.83 94.43 97.32 70.15 74.23 65.63
Ours 91.82 94.63 67.15 71.07 67.32 97.28 98.32 73.52 77.84 64.82 94.93 97.85 70.62 75.14 66.23
We extend and evaluate our method on the medical domain, specifically for ICD coding. The results in Tables III and III demonstrate that our proposed loss function consistently surpasses baselines across all metrics in a comprehensive evaluation, considering: (i) Diverse data distribution: full setting (long-tailed distribution) and top-50 frequent labels setting; (ii) Model architectures: RoBERTa, LLaMA, and domain-specialized PLM-ICD; and (iii) ICD code versions: ICD-9 and ICD-10. The consistent performance improvements observed across these multidimensional evaluation criteria provide substantial empirical evidence for the efficacy and generalizability of our proposed approach.
In the full setting, macro-F1 performance exhibits considerably lower compared to micro-F1, whereas the top-50 setting achieves approximately equal macro and micro-F1 scores. This disparity indicates that extreme long-tailed distributions remain challenging for both the MSCL framework and our method, despite the improvements achieved.
Table III reports that our method achieves superior results on MIMIC-IV-ICD9-Full compared to MIMIC-III-Full, despite both datasets employing identical ICD-9 coding standards. This marked performance differential can be attributed primarily to the more extensive training corpus available in MIMIC-IV-ICD9-Full (see in Table I). While MIMIC-IV-ICD10-Full similarly comprises a substantial volume of clinical data, its considerably expanded label taxonomy introduces increased representational sparsity and presents additional computational and methodological challenges [18]. Moreover, the MIMIC-IV-ICD10-50 dataset demonstrates consistent performance metrics in this restricted setting, providing empirical evidence that label space dimensionality constitutes a critical determinant of model training efficacy.
Comparative analysis of model performance reveals that Llama significantly outperforms RoBERTa across evaluation metrics, a finding attributable to scaling laws of LLMs and the extensive knowledge and training corpus during the pre-training phase [28, 29]. Although LLMs demonstrate considerable efficacy in domain-specific applications [30], our results indicate that PLM-ICD consistently surpasses both RoBERTa and Llama across all experimental configurations. This hierarchical performance pattern aligns with theoretical expectations, as PLM-ICD incorporates architecture and training paradigms specifically optimized for automated ICD coding tasks [24]. Despite the increasing generalization capabilities of foundation models in diverse applications, significant questions persist regarding their capacity to achieve state-of-the-art performance on highly specialized tasks, particularly within the medical domain, without substantial domain-specific training or parameter-efficient adaptation techniques [31]. Contemporary research on foundation model applications in biomedical domain has predominantly relied on specialized adaptation methods tailored to specific domain requirements. The comparative advantages of domain-specific pre-training becomes particularly evident following the development of initial foundation model architectures, as exemplified by widely implemented medical models such as Med-PaLM [32] and Med-Gemini[31].
Therefore, compared with the enhancements via the contrastive training phase, the intrinsic knowledge within LLMs contributes substantially more to ICD coding efficacy. In particular, domain-specific knowledge representations emerge as critical factors of LLMs performance in medical applications.
V Related Work
Contrastive learning aims to learn a representation of data such that similar instances are close together in the representation space, while dissimilar instances are far apart. Compared to self-supervised contrastive learning, such as SimCLR [7] and MoCo [8], Khosla et al. [5] proposed supervised contrastive learning, which fully leverages class annotation information to enhance representations within the contrastive learning framework. Recent studies have extended supervised contrastive learning from single-label to multi-label scenarios by exploiting the additional information inherent in multi-label tasks. Zhang et al. [33] proposed a hierarchical multi-label representation learning framework specifically designed to utilize comprehensive label information while preserving hierarchical inter-class relationships.
In subsequent research, Zhang and Wu [6] developed Multi-Label Supervised Contrastive Learning (MulSupCon), featuring a novel contrastive objective function that expands the positive sample set based on label overlap proportions. Similarly, the Jaccard Similarity Probability Contrastive Loss (JSPCL) [34] employed the Jaccard coefficient [35] to calculate label similarity between instances, sharing conceptual foundations with MulSupCon [6] and MSC loss [19] that those approaches primarily focus on similarity only, but ignoring dissimilarity.
Despite these advancements, the intricate relationships and dependencies between multi-label samples have yet to be fully elucidated. To address this gap, we introduce multi-label relations and formalize the concepts of similarity and dissimilarity. Inspired by the idea of re-weighting of logit adjustment [36], focal loss [26] and class-balanced loss [37], we leverage the similarity and dissimilarity factors to re-weight the contrastive loss, thereby enhancing discriminative power in multi-label scenarios.
VI Conclusion
Multi-label classification poses a compelling challenge in applying contrastive learning due to the diverse ways of defining relations between multi-label samples. In this paper, we introduce multi-label relations and formalize the concepts of similarity and dissimilarity. Then, we propose a Similarity-Dissimilarity Loss for MSCL, which dynamically re-weights the loss by the combination of similarity and dissimilarity factors. We provide theoretical grounded proof for our method through rigorous mathematical analysis that supports the formulation and effectiveness of the proposed loss function. Then, We conduct a comprehensive experiments, considering data modality, domain-specific, data distribution and backbone models to further evaluation our method. The results confirm the effectiveness and robustness of our method in image, text and medical domain.